TheSmartOne (thesmartone):

How would we solve this: \[\lim_{x \rightarrow 0} \frac{\sinh x - \sin x}{x^{3}}\] Without using a graphing tool.

OpenStudy (unavailabilityy):

._.

zepdrix (zepdrix):

Our limit is giving us the indeterminate form 0/0. So I guess we could apply L'Hopital's Rule. Have you learned about this yet?

TheSmartOne (thesmartone):

Nope I noticed it was giving 0/0 also but I thought about rationalizing the numerator, I don't think we can do that though

zepdrix (zepdrix):

Hmm :\

TheSmartOne (thesmartone):

yeah, which is why I'm confused as well xD We can break apart the fraction and get \[\lim_{x \rightarrow 0} \frac{\sinh x - \sin x}{x^{3}}\] \[\lim_{x \rightarrow 0}( \frac{\sinh x}{x^3} - \frac{\sin x}{x^{3}})\] and then \[\lim_{x \rightarrow 0} \frac{\sinh x}{x^3} -\lim_{x \rightarrow 0} \frac{\sin x}{x^{3}}\] \[\lim_{x \rightarrow 0} \frac{\sinh x}{x^3} -\lim_{x \rightarrow 0} (\frac{\sin x}{x}\cdot \frac{1}{x^2})\] mhmmmm

zepdrix (zepdrix):

Ya I can't seem to figure out the algebraic approach for this one... I know L'Hopital will work. You have to apply it three times though.\[\large\rm ~~~\lim_{x\to0}\frac{\sinh x-\sin x}{x^3}\quad=\quad \lim_{x\to0}\frac{\cosh x-\cos x}{3x^2}\]\[\large\rm =\lim_{x\to0}\frac{\sinh x+\sin x}{6x}\quad=\quad \lim_{x\to0}\frac{\cosh x+\cos x}{6}\] Differentiating the numerator and denominator separately. Checking to make sure you still have your indeterminate form 0/0 after each application. Then finally you can plug in x=0 once you've gotten rid of the x in the denominator. Leading to 2/6 or 1/3.

TheSmartOne (thesmartone):

L'Hopital's rule is that easy? ;o woah, thanks zeppy o;